Calculate the no. of electrons in 62 gm nitrate ion.

Grade:10

1 Answers

Bhavya
7 years ago
Dear student
Nitrate ion = NO3-
Molecular mass of nitrate = 14 + 3*16 = 62 g
No of moles in 62 g of nitrate thus would be = 1 mole
Now we know that 1 mole contains 6.022 * 1023molecules
1 molecule contains = 32 electrons
therefore,
6.022 * 1023 molecules will contain = 32 * 6.022 * 1023electrons